OpenStudy (anonymous):

What is y^4-y^3=0 How do you factor this out?

Nnesha (nnesha):

what's the common factor ? well its soo easy to find common factor of variables `variable with the smallest degree would be the common factor `

OpenStudy (anonymous):

How would you do it though?

OpenStudy (danjs):

you are really multiplying by y^2 / y^2 this will pull a y^2 term out of each of those

Nnesha (nnesha):

we should take out the common factor so first ) what's a common factor ???

OpenStudy (danjs):

left with y^2*(each divided by y^2)

OpenStudy (anonymous):

@DanJS my book did it by y(y^3-1) =y(y-1)(y^2+y+1) I dont understand how they do it

Nnesha (nnesha):

well ^^take out y from y^4 -y^3 \[\huge\rm y(y^3-1)\] y^3-1 is same as (y^3 -1^3)

Nnesha (nnesha):

then you can apply the rule \[\huge\rm x^3-y^3=(x-y)(x^2+xy+y^2)\]

OpenStudy (danjs):

is the original problem y^4 - y^3

OpenStudy (danjs):

= y^3 * (y-1) = 0

OpenStudy (danjs):

y is 0 or 1

OpenStudy (anonymous):

Yes

Nnesha (nnesha):

i guess it's y(y^3-1) ??

OpenStudy (anonymous):

Shouldn't it factor out as y^3(y-1)?

OpenStudy (anonymous):

Doesn't y(y^3-1) = y^4-y

Nnesha (nnesha):

ye....... and then set it equal to zero if you want to solve to for y but factor is y^3(y-1) that looks correct 2 me :/

OpenStudy (mathstudent55):